find the answer for the following question a wheel is initially at rest is subjected to an angular acceleration and it completes 10 rotations in time 't', then the time taken by it to complete next 10 rotations

Grade:12

1 Answers

Vikas TU
3 years ago
θ = 1/2 a t²
For 10 rotations :
10 × 2π = 1/2 a t²
a = 40 π/t² - - - - - - - - ( i )
As total rotations = 20, with the help of eq i
T² = 80 π/a
=2t²
T = √2 t
After 10 rotations = ( √2 - 1 ) t